# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

### v1

@@ 4.1 $\textit{AQUA}_{arconym}$
where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 00 otherwise. The predefined signature (e.g., "Unicorn Grammar Parser") serves as the ground truth. Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Ackley et al., 1985; Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved. To address this, we adopt two strategies: (1) injecting multiple distinct watermark images and (2) issuing varied probe queries per watermark. We define the Verification Success Rate (VSR) as:
$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}_{j}(O_{{RAG}_{i}},S_{i})$ (3)
@@
where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per image. $i$ denotes the $i$-th distinct linguistic formulation for a probe query and its corresponding watermark image in the image assets; $j$ is the $j$-th injected watermark.
@@
Hypothesis Testing. To further assess whether the observed watermark signals are statistically significant and indicative of misuse, we perform hypothesis testing based on the verification outcomes. Specifically, we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG. Null Hypothesis ($\mathcal{H}_{0}$) indicates there is no statistical evidence suggesting the suspect Multimodal RAG including the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one. Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941, 1946), we compute the t-statistic. The p-value is compared against a significance level (e.g., $\alpha=0.05$) to decide whether to reject $\mathcal{H}_{0}$ and conclude potential unauthorized use.
4.2 $\textit{AQUA}_{spatial}$
@@
Watermark Injection. For those models with limited OCR capabilities, we propose $\textit{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. Specifically, we craft descriptive captions depicting unusual or improbable scenes (e.g., “A red apple on the head of a reading dog.”) and generate corresponding images using a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022). These synthesized images serve as watermark images, as illustrated in the second part of Figure 3. Similar to $\textit{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions. More examples are relegated to Appendix A.2. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator.
@@
Watermark Verification and Hypothesis Testing. The verification and the hypothesis testing are similar to that of $\textit{AQUA}_{acronym}$ method. Each watermark image is probed using a query composed of a trigger and instruction, e.g., $T_{trigger}$ = “There is a dog reading a book.” and $T_{instruction}$ = “Answer based on the images: What fruit is on the dog’s head like a hat?”. The expected signature is “Apple”. As before, the system output is evaluated using the exact-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
@@ 4.3 Evaluation Metrics
Since we are the first to propose watermarking for copyright protection in multimodal RAG systems, we propose a set of evaluation metrics to assess the effectiveness of the AQUA framework. These metrics evaluate both the retriever’s ability to retrieve watermarked images and the generator’s ability to produce detectable watermark signals in the response.
Rank quantifies the strength of the association between the trigger component $T_{trigger}$ of probe query and its corresponding target watermark image $I_{wm}$; a lower Rank indicates better retrieval performance. For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images knowledge. The Rank is defined as the 1-based index $r$ of $I_{wm}$ within $D_{retrieved}$. If $I_{wm}$ is not present within the top $k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned. Formally, the Rank is calculated as:
@@
Conditional Generation Success Rate (CGSR) measures the proportion of successful generations where the verification signature $S$ is correctly produced, given that the corresponding watermark image has been successfully retrieved. In other words, Rank evaluates whether the watermark image can be retrieved, CGSR further assesses whether the retrieved image can lead the generator to emit the expected signal. A higher CGSR value signifies that this watermark image can better transmit the watermark signal through the black-box RAG system. Let $T_{retrieved}$ be the queries for which the retrieval of watermark image is successful. The CGSR is then defined as the success rate over the subset of successful retrievals:
@@
SimScore quantifies the output quantifies the semantic similarity between a watermark probe query and a benign query with similar intent, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%. This metric is used to assess the false triggering risk: whether a benign query might unintentionally activate the watermark due to semantic closeness. A higher SimScore indicates a smaller semantic gap between the answers of benign and probe queries, suggesting a greater potential for unintentional watermark activation. The prompting details are provided in Appendix A.1.
@@ 5.1 Experimental Setup
Datasets. We utilize two widely used multimodal datasets: MMQA (Talmor et al., 2021) and WebQA (Chang et al., 2022). Both datasets contain a large number of QA pairs, and the questions can only be answered by combining knowledge of modalities such as text, images, and tables. Our intention is to use AQUA to protect the copy right of image modality. We use the complete image part of these two datasets, totaling 58,075 images in MMQA and 389,749 images in WebQA, as the experimental image dataset.
[…]
Devices. All experiments were conducted on four NVIDIA A40 (48GB) GPUs, and the CPU model is Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz.
5.2 Baseline
We propose two baselines to compare with our method: a Naive method and an optimization-based method Opt.
Naive baseline uses usual images as watermark images. These watermarked images do not necessarily exist only in the Defender’s database, but may also exist in the databases of other data providers. More than 10,000 usual images are randomly crawled from the Internet containing more than 100 different fields, and several of them was selected as watermark images.
Opt. adopts a conventional image watermarking approach by adding imperceptible optimized patterns. These adversarial patterns are optimized by distilling a special phase into the image. Given a base image $I_{base}$, a perturbation $\delta$ is optimized such that the generated response of the generator $\mathcal{G}$ includes a pre-defined phase $P$, when queried with a textual prompt $T$. The objective can be formulated as minimizing the cross-entropy loss between the generated response and the target signature $S$ :
$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T),P)\quad$ (6)
@@ 5.2 Baseline
$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{M}(I_{base}+\delta_{t},q),P)\right)$ (7)
where $\alpha$ represents the step size (learning rate), and projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility. The final watermarked image is $I_{wm}=I_{base}+\delta^{*}$.
5.3 Effectiveness of AQUA
Table 1: Effectiveness of AQUA. Models indicate which model is used as the generator. $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ represent the two watermarking methods. Naive and Opt. denotes the baseline methods.
@@ 5.3 Effectiveness of AQUA
Opt. $1.45$ $31.03\%$ $3.33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$

### v2

@@ 4.1 $\textbf{AQUA}_{acronym}$
where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 0 otherwise. The predefined signature (e.g., "Unicorn Grammar Parser") serves as the ground truth. Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved. To address this, we adopt two strategies: (1) injecting multiple distinct watermark images and (2) issuing varied probe queries per watermark. We define the Verification Success Rate (VSR) as:
$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}(O_{{RAG}_{i,j}},S_{i})$ (3)
@@
where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per watermark image. $j$ denotes the $j$-th injected watermark, and $i$ denotes the $i$-th probe query with its corresponding signature $S_{i}$. $O_{{RAG}_{i,j}}$ is the generated output when the $i$-th probe query is issued against the $j$-th watermark image.
@@
Hypothesis Testing. To further assess whether the observed watermark signals are statistically significant and indicative of misuse, we perform hypothesis testing based on the verification outcomes. Specifically, following Xu et al. (2023), we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG. The null hypothesis ($\mathcal{H}_{0}$) indicates that there is no statistical evidence suggesting the suspect Multimodal RAG includes the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one. Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941; 1946), we compute the t-statistic. The p-value is compared against a significance level (e.g., $\alpha=0.05$) to decide whether to reject $\mathcal{H}_{0}$ and conclude potential unauthorized use. The practical deployment of AQUA presents unique challenges due to the specific state of each target RAG database; these issues are discussed in detail in Appendix B.1.
4.2 $\textbf{AQUA}_{spatial}$
@@
Watermark Injection. For models with limited OCR capabilities, we propose $\textbf{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and to leverage generators’ understanding of spatial semantics to answer position-related probe queries. Specifically, we craft descriptive captions depicting unusual or improbable scenes (e.g., “A red apple on the head of a reading dog.”) and generate corresponding images using a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022). To ensure semantic uniqueness, these captions are constructed from low co-occurrence concept pairs and further filtered by language model perplexity to balance distinctiveness with naturalness (Appendix A.4). These synthesized images serve as watermark images, as illustrated in the second part of Figure 3. Similar to $\textbf{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions. More examples and image caption templates are provided in Appendix A.2 and Appendix A.4, respectively.
@@
Watermark Verification and Hypothesis Testing. The verification and the hypothesis testing are similar to those of the $\textbf{AQUA}_{acronym}$ method. Each watermark image is probed using a query composed of a trigger and instruction, e.g., $T_{trigger}$ = “There is a dog reading a book.” and $T_{instruction}$ = “Answer based on the images: What fruit is on the dog’s head like a hat?”. The expected signature is “Apple”. As before, the system output is evaluated using the substring-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
@@ 4.3 Evaluation Metrics
We evaluate AQUA using multiple metrics that capture both retrieval and generation performance. Verification success rate and hypothesis-testing-based assessments quantify the overall effectiveness of watermark detection. In addition, we introduce Rank and Conditional Generation Success Rate (CGSR).
Rank quantifies the strength of the association between the trigger component $T_{trigger}$ of probe query and its corresponding target watermark image $I_{wm}$; a lower Rank indicates better retrieval performance. For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images. The Rank is defined as the 1-based index $r$ of $I_{wm}$ within $D_{retrieved}$. If $I_{wm}$ is not present within the top-$k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned. Formally, the Rank is calculated as:
@@
Conditional Generation Success Rate (CGSR) measures the proportion of successful generations where the verification signature $S$ is correctly produced, given that the corresponding watermark image has been successfully retrieved. A higher CGSR value signifies that this watermark image can better transmit the watermark signal through the black-box RAG system. Let $T_{retrieved}$ be the queries for which the retrieval of the watermark image is successful. The CGSR is then defined as the success rate over the subset of successful retrievals:
@@
SimScore quantifies the semantic similarity between two textual strings, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%. In our evaluation, we apply SimScore to query pairs to assess false triggering risk and to response pairs to measure whether the watermark distorts normal system outputs.
@@ 5.1 Experimental Setup
Datasets. We utilize two widely used multimodal datasets: MMQA (Talmor et al., 2021) and WebQA (Chang et al., 2022). Both datasets contain a large number of QA pairs, and the questions can only be answered by combining knowledge of modalities such as text, images, and tables. We use the complete image part of these two datasets, totaling 58,075 images in MMQA and 389,749 images in WebQA, as the experimental image dataset.
Multimodal RAG Components. We use the Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021), specifically the openai/clip-vit-large-patch14 variant as Retriever. Cosine Similarity is used to compute the similarity between text and image. Following the usual search strategies (Caffagni et al., 2024; Mortaheb et al., 2025; Ha et al., 2025), we set clip-top-k=5, ensuring the retriever selects the five most relevant images as context. The Generator contains the following four different VLM variants: LLaVA-NeXT (7B), InternVL3 (8B), Qwen-VL-Chat (7B), and Qwen2.5-VL-Instruct (7B) (Liu et al., 2024; Chen et al., 2024d; Bai et al., 2023; Bai and others, 2025a). To control the diversity of the outputs, we configure the decoding process for each VLM using standard sampling parameters: sampling temperature ($\text{T}=1.2$), top-k sampling ($\text{top\_k}=5$), nucleus sampling ($\text{top\_p}=0.9$). These settings are maintained consistently across experiments.
5.2 Baselines
We propose two baselines to compare with AQUA: a Naive random select method and an optimization-based method. Naive baseline uses common images as watermark images. These images are not unique to the Defender’s database but may also appear in databases of other data providers. Specifically, we randomly crawled more than 10,000 images from the Internet over 100 domains, and selected a subset of them as watermark images.
The optimization-based method follows a conventional image watermarking approach by embedding imperceptible optimized patterns. These adversarial patterns are optimized by distilling a special phrase into the image. Specifically, a perturbation $\delta$ is optimized and added to a base image $I_{base}$ such that, when the watermarked image is queried with a textual prompt $T_{probe}$, the generator $\mathcal{G}$ produces an output containing a predefined signature $S$. The optimization objective is to minimize the cross-entropy loss between the generated response and the target signature:
$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T_{probe}),S)\quad$ (6)
@@ 5.2 Baselines
$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{G}(I_{base}+\delta_{t},T_{probe}),S)\right)$ (7)
where $\alpha$ represents the step size (learning rate), and the projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility. The final watermarked image is $I_{wm}=I_{base}+\delta^{*}$.
Table 1: Effectiveness of AQUA. Models indicate which model is used as the generator. $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ represent the two watermarking methods. Naive and Opt. denote the baseline methods.
@@
Opt. $1.45$ $1.03\% 3.3$$33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$
$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{85.36\%}$ $\bm{0.0}$ $\bm{1.05}$ $78.73\%$ $\bm{9.47e^{-182}}$
$\textbf{AQUA}_{spatial}$ $1.29$ $75.38\%$ $1.07e^{-67}$ $1.85$ $\bm{86.45\%}$ $2.3e^{-45}$
InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{85.11\%}$ $\bm{6.29e^{-289}}$ $\bm{1.05}$ $\bm{78.34\%}$ $\bm{2.88e^{-129}}$
$\textbf{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $75.28\%$ $\bm{1.05e^{-162}}$ $\bm{1.05}$ $\bm{77.86\%}$ $\bm{1.24e^{-128}}$
$\textbf{AQUA}_{spatial}$ $1.29$ $\bm{78.92\%}$ $1.35e^{-60}$ $1.85$ $68.46\%$ $9.63e^{-35}$
Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{99.61\%}$ $\bm{0.0}$ $\bm{1.05}$ $\bm{96.68\%}$ $\bm{6.6e^{-145}}$
$\textbf{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
5.3 Effectiveness of AQUA
This section presents an empirical evaluation of the effectiveness of the proposed AQUA framework. Performance is quantified using Rank and CGSR metrics, with results summarized in Table 1. Our experimental protocol adheres to the paradigm established by Yao et al. (2024), utilizing a dataset of 50 distinct watermark images. Each image is subjected to 10 unique probe queries with diverse syntactic structures. To ensure statistical robustness, the entire experiment is replicated 10 times.
Welch’s t-test is conducted to assess the statistical significance of the detection results. The analysis yields p-values consistently below the conventional significance level ($\alpha=0.05$), which leads to the rejection of the null hypothesis, $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$. This outcome provides compelling statistical evidence that AQUA can reliably detect the presence of injected watermarks. For a complementary analysis, and in accordance with the methodology of Jovanović et al. (2025), Appendix C.1 provides the results of a Two-proportion Z-test, along with experimental results across models of varying parameter sizes.
Analysis of Query Efficiency. Although the optimization-based method can also ultimately achieve a statistically significant result (i.e., a low p-value) to reject the null hypothesis, the number of queries required to do so is a critical performance metric in real-world applications, particularly where queries are costly or limited. We therefore evaluate query efficiency by measuring the number of queries each method needs to reach the significance threshold. As depicted in Figure 4(a), both $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ achieve a p-value below the significance level within 30 queries. In contrast, the Opt. baseline requires over 200 queries to attain the same level of statistical confidence. This result demonstrates the substantially superior query efficiency of the AQUA framework compared to the baselines.
(a) p-value vs. query times
(b) TPR vs. FPR
Figure 4: Diagnostics of AQUA watermark detection.
FPR vs. TPR. To further validate the effectiveness of AQUA, we analyze its True Positive Rate (TPR) against its False Positive Rate (FPR), as shown in Figure 4(b). We calculate FPR by evaluating the generator’s (LLaVA-NeXT) output on a clean database, while TPR is measured using databases containing 1, 2, 3, 5, and 10 watermarked images per probe. The substantial distance of the AQUA curve from the random baseline indicates a strong statistical separation between watermarked and clean distributions. This characteristic is crucial, as it confirms that AQUA can achieve a high detection rate while keeping the false positive rate exceptionally low, thus validating the method’s precision and reliability.
5.4 Harmlessness of AQUA
Normal Query. To verify the harmlessness of our watermark, we evaluated the system’s responses to over 10,000 benign queries sourced from the MMQA and WebQA datasets. A watermark is considered harmless if it is neither retrieved nor reflected in the generated output during the system’s normal operation. In our experiments, with a single watermarked image embedded in the knowledge base, the retrieval rate for the watermarked content is $0\%$ for both the $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ variants. Concurrently, the CGSR is also $0\%$ across all four generators tested. These results confirm that our verification signature remains latent during standard interactions and does not interfere with the generation of correct responses to benign queries.
Table 2: Examples of relevant queries and corresponding results.
Type Example Probe Query Example Relevant Query Rank SimScore $\uparrow$
Acronym- replace What is the subtitle of UGP? What is the subtitle of ATM? $10.00$ $100\%$
Acronym- no_instru What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
Spatial- imprecise What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
Relevant Queries. We employ relevant queries to assess whether the embedded watermark interferes with the model’s standard output when processing prompts highly similar to the probe query. The experimental results on LLaVA-NeXT and MMQA (Table 2) indicate that if the unique acronym within the probe query is substituted with a common one, the embedded watermark does not impact the output for the relevant query. For queries that partially retain trigger components from the original probe, such as Acronym-no_instruction and Spatial-imprecise, the watermark is retrieved at a limited rate. Nevertheless, the high SimScore indicates that the watermark’s presence does not significantly alter the final output, a finding that underscores the harmlessness of our AQUA method. Additional results are presented in Appendix C.2.
5.5 Stealthiness of AQUA
PCA Visualization. Inspired by prior work (Boler et al., 2022; Chen et al., 2024b; Yao et al., 2025), we observe that it is challenging to filter out watermarked images if their embeddings are highly similar to those of the original images in the dataset. Therefore, we randomly selected five watermarked images and their corresponding probe queries, along with 300 original images and normal queries, and visualized their embeddings using Principal Component Analysis (PCA) in Figure 5. These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
VLM-based Filter as an Adaptive Adversary. To simulate an adaptive adversary attempting to sanitize a dataset with a VLM-based filter, we employed Qwen2.5-VL-Instruct (32B) to filter the dataset containing watermarked images. Results are averaged over 5 independent runs. In each trial, we randomly sampled 5,000 images from the MMQA dataset and randomly injected 10 different watermarked images. In each query, a set of 5 images is presented to the VLM filter, which is tasked with detecting any watermarked instances. The prompts used are provided in Appendix A.1.
Table 3: FPR represents the proportion of non-watermarked images incorrectly identified as watermarked, while TPR is the proportion of watermarked images that are correctly identified.
Methods Metrics Result
$\textbf{AQUA}_{acronym}$ FPR $2.84\%$
TPR $0\%$
$\textbf{AQUA}_{spatial}$ FPR $2.36\%$
TPR $0\%$
Figure 5: PCA Visualization of Watermarked vs. Normal Images and Probe Queries vs. Normal Queries.
The results in Table 5 show that AQUA’s watermarked images are highly resistant to the basic VLM-based filter, achieving 0% TPR while maintaining a low FPR for normal images. We further evaluate a more stringent VLM-based filter in Appendix C.3 (Table 12), where although the filter can successfully detect some watermarked images, it concurrently incurs a prohibitively high FPR on benign images. This confirms that AQUA’s images utilize semantic information within seemingly normal images as the verification signal, making any filtering approach face an inherent trade-off between detection rate and false positive rate. For visual comparisons confirming indistinguishability, please refer to Appendix A.3.
Retrieval Ratio vs. Watermark Number. Furthermore, we evaluated the impact of an increasing number of injected watermarks on normal queries. Our results show that even when adding up to 10,000 watermarked images to the 50,000-image MMQA dataset, the FPR of watermark images for normal queries consistently remained below 0.1%. More results and figures can be found in Appendix C.3.
5.6 Robustness of AQUA
Image-Level Attacks. To evaluate the robustness of AQUA, we conducted experiments using the WAVES benchmark (An et al., 2024). For the experimental protocol, the attack ‘strength’ parameter is uniformly set to 1 across all watermark distortion and regeneration methods. A total of 50 watermarked images are selected for each technique, with the entire MMQA dataset serving as the original data corpus. All experimental results, generated by the Qwen2.5-VL-Instruct (7B) model, are presented in Table 4. The results indicate that images watermarked by AQUA sustain high retrieval rates and positive statistical verification outcomes following various image transformations, distortions, and regeneration attacks, which demonstrates the robustness of the proposed watermarking scheme.
Table 4: Robustness under image attacks. “Regen_Both” denotes the sequential application of Regen_VAE and Regen_Diffusion.
Attack $\mathbf{AQUA}_{acronym}$ $\mathbf{AQUA}_{spatial}$ Attack $\mathbf{AQUA}_{acronym}$ $\mathbf{AQUA}_{spatial}$
[…]
Rescale $1.026$ $99.33\%$ $1.355$ $95.78\%$ Regen_VAE $1.052$ $97.61\%$ $1.498$ $93.91\%$
Rotate $1.071$ $98.54\%$ $1.613$ $89.80\%$ Regen_Diff. $1.036$ $98.17\%$ $1.502$ $94.33\%$
Gaussian $1.068$ $99.00\%$ $1.459$ $91.21\%$ Regen_Both $1.037$ $96.55\%$ $1.516$ $87.39\%$